OpenStudy (anonymous):

For the inequality x+2 over x+6>0. Need clarification on interval notation and algebraic notation solution

OpenStudy (anonymous):

for algebraic notation I have x>-2 or x<-6. Anyone know if this is correct and what the interval notation looks like?

OpenStudy (anonymous):

This is my last problem on my online assignment and I can't get any more wrong or i have to start over so i'm trying to make sure I get the right format. In case you're wondering why I have this up here again. ha

OpenStudy (anonymous):

There is this thing called google. I used it to find this particular website. You might wanna check it out. http://zonalandeducation.com/mmts/miscellaneousMath/intervalNotation/intervalNotation.html

OpenStudy (anonymous):

yeah, I looked at that too. I just don't see how this looks different than the algebraic answer.

OpenStudy (anonymous):

um, algebraic notation is x>-2 or x<-6 the interval notation is x = { (- infinity, -6) , (-2, infinity)}

OpenStudy (anonymous):

how do the two look same to you?

OpenStudy (anonymous):

do they have to have the {} around them? So for x+1/x+7>0 here I got alg: x>-1 or x<-7. will the notation one be \[\left\{ -\infty,-7 \right\}\left\{ -1,\infty \right\}\]

OpenStudy (anonymous):

I just posted a link as to how interval notation works. Read it.

OpenStudy (anonymous):

It is better than reading my explanation.

OpenStudy (anonymous):

x = (5,7) means that x lies between, BUT NOT including 5 and 7. x = [5,7] means that x lies between AND including 5 and 7 x = [5,7) means that x lies between 5 and 7, including 5, but not 7.

OpenStudy (anonymous):

Yeah I was looking at that. Just wasn't positive on which one to put for this problem. Wasn't even sure if I have the right solution to go off of

OpenStudy (anonymous):

okay let me ask you a question what does x>5 mean to you?

OpenStudy (anonymous):

x=-5?

OpenStudy (anonymous):

wait. nvm

OpenStudy (anonymous):

what? NO! how does that even make any sense? Just tell me in plain english what you mean by x > 5.

OpenStudy (anonymous):

x is greater than 5.. I didn't realize how simple the question was. I misread what you wrote first.

OpenStudy (anonymous):

yes. So x is greater than 5. does that mean that the value of x lies between 5 and infinity?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, but does it also mean that x could be 5?

OpenStudy (anonymous):

no because it isn't greater or equal to

OpenStudy (anonymous):

okay good. so x lies between 5 and infinity, including infinity, BUT NOT including 5. so we write x = (5, infinity]

OpenStudy (anonymous):

just remember, parentheses ( ) mean not including. square brackets [ ] mean including

OpenStudy (anonymous):

\[[-\infty,-7)(-1,\infty]\] Sorry for being so annoying I just want to make sure this one looks correct before I submit it. Think this is the answer to the x+1/x+7>0?

OpenStudy (anonymous):

Also, thanks for explaining that. It makes sense now. Just want to get the correct format

OpenStudy (anonymous):

yes, that is right. that should be the answer to the second one. you are welcome. it is better to understand what you are talking about rather than just follow what others say.

OpenStudy (anonymous):

it wont et me only put the [ on one side. maybe it needs to be in this format.. [-inf(-7-1)int]

OpenStudy (anonymous):

no, that is incorrect format. you dont need to put square brackets for infinity, because infinity can never really be reached. so it doesn't matter.

OpenStudy (anonymous):

you should have read the article I posted. that link explains what you need to do. Well, let us get just a bit more complicated. Using interval notation we will show the set of number that includes all real numbers except 5. First, stated as inequalities this group looks like this: The statement using the inequalities above joined by the word or means that x is a number in the set we just described, and that you will find that number somewhere less than 5 or somewhere greater than 5 on the number line. In interval notation a logically equivalent statement does not use the word or, but rather a symbol for what is called the union of two groups of numbers. The symbol for union coincidentally looks like a U, the first letter of union. However, it is really not a letter of the alphabet. Here is what the union symbol looks like: U So, the group of numbers that includes all values less than 5 and all values greater than 5, but does not include 5 itself, expressed as interval notation looks like this: \[(-\infty,5) \cup (5,+\infty)\]